40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 25 17:42:11 2021, Anchal, Paco, Summary, ALS, ALS Single Arm Noise Budget ALS_IR_b.svgALS_Single_Arm_IR.pdf
    Reply  Thu May 27 11:03:15 2021, Anchal, Paco, Summary, ALS, ALS Single Arm Noise Budget ALS_Single_X_Arm_IR.pdfALS_OOL_with_Ref.pdf
       Reply  Fri May 28 17:32:48 2021, Anchal, Summary, ALS, Single Arm Actuation Calibration with IR ALS Beat SingleArmActCalwithIRALSBeat.pdfstateSpaceModel.zip
          Reply  Tue Jun 1 16:55:32 2021, Anchal, Paco, Summary, ALS, Single Arm Actuation Calibration with IR ALS Beat SingleArmActCalwithIRALSBeat-1306624785.pdf
             Reply  Tue Jun 8 11:40:53 2021, Anchal, Paco, Summary, ALS, Single Arm Actuation Calibration with IR ALS Beat ITMX_Cal_Noise_Spectrum_1307143423.pdf
                Reply  Wed Jun 9 18:29:13 2021, Anchal, Paco, Summary, ALS, Check for saturation in ITMX SOS Driver 
                Reply  Fri Jul 9 15:39:08 2021, Anchal, Summary, ALS, Single Arm Actuation Calibration with IR ALS Beat [Correction] ITMX_calibration_With_ALS_Beat.pdf
Message ID: 16161     Entry time: Tue May 25 17:42:11 2021     Reply to this: 16164
Author: Anchal, Paco 
Type: Summary 
Category: ALS 
Subject: ALS Single Arm Noise Budget 

Here is our first attempt at a single-arm noise budget for ALS.

Attachment 1 shows the loop diagram we used to calculate the contribution of different noises.

Attachment 2 shows the measured noise at C1:ALS-BEATX_PHASE_FINE_OUT_HZ when XARM was locked to the main laser and Xend Green laser was locked to XARM.

  • The brown curve shows the measured noise.
  • The black curve shows total estimated noise from various noise sources (some of these sources have not been plotted as their contribution falls off the plotting y-lims.)
  • The residual frequency noise of Xend green laser (AUX) is measured by measuring the PDH error monitor spectrum from C1:ALS-X_ERR_MON_OUT_DQ. This measurement was converted into units of V by multiplying it by 6.285e-4 V/cts. This factor was measured by sending a 43 Hz 100 mV sine wave at the readout point and measuring the output in the channel.
  • This error signal is referred to AUX_Freq input in the loop diagram (see attachment 1) and then injected from there.
  • All measurements were taken to Res_Disp port in the 'Out-of-Loop Beat Note' block (see attachment 1).
  • In this measurement, we did not DAC noise that gets added when ALS loop is closed.
  • We added ADC noise from Kiwamu's ALS paper after referring it to DFD input. DFD noise is also taken from Kiwamu's ALS paper data.

Inference:

  • Something is wrong above 200 Hz for the inclusion of AUX residual displacement noise. It is coming off as higher than the direct measured residual noise, so something is wrong with our loop diagram. But I'm not sure what.
  • There is a lot of unaccounted noise everywhere from 1 Hz to 200 Hz.
  • Rana said noise budget below 1 Hz is level 9 stuff while we are at level 2, so I'll just assume the excess noise below 1 Hz is level 9 stuff.
  • We did include seismic noise taken from 40m noise budget in 40m/pygwinc. But it seems to affect below the plotted ylims. I'm not sure if that is correct either.

Unrelated questions:

  • There is a slow servo feeding back to Green Laser's crystal temperature by integrating PZT out signal. This is OFF right now. Should we keep it on?
  • The green laser lock is very unreliable and it unlocks soon after any signal is being fed back to the ETMX position.
  • This means, keeping both IR and green light locked in XARM is hard and simultaneous oscillation does not last longer than 10s of seconds. Why is it like this?
  • We notice that multiple higher-order modes from the green laser reach the arm cavity. The HOMs are powerful enough that PDH locks to them as well and we toggle the shutter to come to TEM00 mode. These HOMs must be degrading the PDH error signal. Should we consider installing PMCs at end tables too?
Attachment 1: ALS_IR_b.svg  337 kB  Uploaded Tue May 25 18:43:38 2021  | Hide | Hide all | Show all
ALS_IR_b.svg
Attachment 2: ALS_Single_Arm_IR.pdf  128 kB  Uploaded Tue May 25 18:53:04 2021  | Show | Hide all | Show all
ELOG V3.1.3-